# Generalized Virial Identities: Radial Constraints for Solitons, Instantons, and Bounces

### v1

@@ 4.3 Nielsen-Olesen vortices
The Nielsen-Olesen vortex NielsenOlesen1973 is the fundamental topological defect in the Abelian Higgs model. For a unit-winding vortex with cylindrical symmetry, the static energy functional reduces to
$E=2\pi v^{2}\int_{0}^{\infty}\left[\frac{1}{2\rho}\dot{a}^{2}+\rho\dot{f}^{2}+\frac{1}{\rho}f^{2}a^{2}+\frac{\beta\rho}{4}(f^{2}-1)^{2}\right]d\rho,$ (46)
@@
The four terms have distinct geometric weights. The gauge kinetic term $\frac{1}{2\rho}\dot{a}^{2}$ arises from the magnetic field $B\propto\dot{a}/\rho$; its $1/\rho$ weight reflects the $1/\rho^{2}$ scaling of magnetic energy density integrated over the azimuthal direction. The scalar kinetic term $\rho\dot{f}^{2}$ represents radial gradient energy with the standard two-dimensional measure. The covariant coupling term $\frac{1}{\rho}f^{2}a^{2}$ arises from the angular covariant derivative $|D_{\theta}\phi|^{2}$; its $1/\rho$ weight reflects the centrifugal nature of angular gradients, with the effective “angular pressure” enhanced at small $\rho$ where the vortex core resides. The Higgs potential term has the standard $\rho$ weight from the measure.
@@ 4.4 Electroweak sphaleron
Unlike the BPST instanton where conformal invariance rendered the $\alpha=1$ identity trivial, the sphaleron’s size and energy are fixed by explicit mass scales: the gauge boson mass $m_{W}$ sets the overall size, while the Higgs mass $m_{H}$ controls the strength of vacuum compression. The interplay between these scales manifests directly in the $\alpha$-dependence of the virial identities.
For the spherically symmetric ansatz with gauge profiles $f_{A}$, $f_{B}$ and Higgs profiles $H$, $K$, the energy functional in the limit $\theta_{W}\to 0$ takes the form MatchevVerner2025:
$E_{\rm sph}$ $=\frac{4\pi}{g^{2}}\int_{0}^{\infty}d\rho\bigg[(\dot{f}_{A})^{2}+(\dot{f}_{B})^{2}+\frac{(f_{A}^{2}+f_{B}^{2}-1)^{2}}{2\rho^{2}}$
$\quad+2m_{W}^{2}\rho^{2}(\dot{H}^{2}+\dot{K}^{2})+m_{W}^{2}U_{\rm int}(f_{A},f_{B},H,K)$
$\quad+\frac{m_{W}^{2}m_{H}^{2}}{2}\rho^{2}(H^{2}+K^{2}-1)^{2}\bigg],$ (52)
where $\rho$ is scaled by $m_{W}$ and the interaction potential is $U_{\rm int}=(f_{A}^{2}+f_{B}^{2})(H^{2}+K^{2})+2(f_{A}H+f_{B}K)^{2}$.
Reading off the geometric weights: gauge kinetic terms have $A_{\rm gauge}=1$, Higgs kinetic terms have $A_{\rm Higgs}=2m_{W}^{2}\rho^{2}$, the centrifugal potential has $B_{\rm cent}=1/(2\rho^{2})$, the interaction potential has $B_{\rm int}=m_{W}^{2}$, and the vacuum potential has $B_{\rm vac}=m_{W}^{2}m_{H}^{2}\rho^{2}/2$. The $\rho^{2}$ factors in the Higgs and vacuum terms reflect the explicit breaking of scale invariance: unlike the conformally invariant Yang-Mills case, here the Higgs mass introduces a definite length scale.
@@ 5 Skyrmions
The Skyrme model provides an effective field theory of pions whose topological solitons, Skyrmions, model baryons in the large-$N_{c}$ limit of QCD Skyrme1962; AdkinsNappiWitten1983. The fundamental $B=1$ Skyrmion has a spherically symmetric hedgehog configuration specified by a radial profile $f(\rho)$ with boundary conditions $f(0)=\pi$ and $f(\infty)=0$.
@@
$E=4\pi\int_{0}^{\infty}\rho^{2}\mathcal{E}(f,\dot{f})\,d\rho,$ (62)
@@ 5.2 Standard Skyrme model
The standard model has $c_{2,0,0}=c_{2,2,0}=1$:
$E=4\pi\int_{0}^{\infty}\left[\rho^{2}\dot{f}^{2}+2\sin^{2}f+2\dot{f}^{2}\sin^{2}f+\frac{\sin^{4}f}{\rho^{2}}\right]d\rho.$ (66)
The four terms represent: Dirichlet (radial) gradient energy $\rho^{2}\dot{f}^{2}$; sigma-model (angular) gradient energy $\sin^{2}f$; Skyrme term $\dot{f}^{2}\sin^{2}f$ mixing radial and angular strains; and centrifugal barrier $\sin^{4}f/\rho^{2}$ that prevents collapse.
@@ 6 Conclusion
The numerical examples illustrate how the $\alpha$-dependence of errors distinguishes between core and tail inaccuracies. For Nielsen-Olesen vortices, the errors grow at negative $\alpha$: the $\alpha=1$ identity held to $0.0005\%$ while $\alpha=-0.5$ showed a $5.7\%$ discrepancy, exposing core-region errors that the global Derrick test averages away. For the Coleman bounce, the opposite pattern emerged: errors grew monotonically with $\alpha$, from $0.001\%$ at $\alpha=-2$ to $0.03\%$ at $\alpha=2$, indicating that numerical inaccuracies were concentrated in the asymptotic tail where the bounce approaches the false vacuum. The $\alpha$-family thus gives a clear signature: core errors manifest at negative $\alpha$, while tail errors manifest at positive $\alpha$.
@@ Acknowledgements.
The author thanks Manuel Labansat for valuable discussions on vacuum decay and bounce solutions.

### v2

@@ 4.3 Nielsen-Olesen vortices
The Nielsen-Olesen vortex NielsenOlesen1973 is the fundamental topological defect in the Abelian Higgs model with Lagrangian
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+|D_{\mu}\phi|^{2}-\frac{\lambda}{4}(|\phi|^{2}-v^{2})^{2},$ (71)
where $D_{\mu}=\partial_{\mu}-ieA_{\mu}$ and $\phi$ is a complex scalar.
For a unit-winding vortex with cylindrical symmetry, the ansatz
$\phi=vf(\rho)e^{i\theta},\qquad A_{\theta}=\frac{1}{e\rho}[1-a(\rho)],$ (72)
reduces the static energy to
$E=2\pi v^{2}\int_{0}^{\infty}\mathcal{G}_{\rho}\,d\rho,\qquad\mathcal{G}_{\rho}=\frac{1}{2\rho}\dot{a}^{2}+\rho\dot{f}^{2}+\frac{1}{\rho}f^{2}a^{2}+\frac{\beta\rho}{4}(f^{2}-1)^{2},$ (73)
@@
The auxiliary quantity for this two-field system is
$\mathcal{C}_{\rho}=\frac{1}{\rho}\dot{a}^{2}+\rho\dot{f}^{2}-\frac{1}{\rho}f^{2}a^{2}-\frac{\beta\rho}{4}(f^{2}-1)^{2}.$ (74)
The gauge kinetic term $\dot{a}^{2}/(2\rho)$ and covariant coupling $f^{2}a^{2}/\rho$ share the weight $\rho^{-1}$, reflecting the magnetic field structure $B\propto\dot{a}/\rho$. The scalar kinetic term $\rho\dot{f}^{2}$ and Higgs potential carry the standard two-dimensional weight $\rho$.
@@
The vortex profiles satisfy $a(0)=1$, $f(0)=0$ (regularity with unit winding) and $a(\infty)=0$, $f(\infty)=1$ (vacuum at infinity)
At the origin ($\rho\to 0$): Regularity requires
$f\sim c_{f}\rho,\quad a\sim 1-c_{a}\rho^{2},\quad\Rightarrow\quad\dot{f}\sim c_{f},\quad\dot{a}\sim-2c_{a}\rho.$ (80)
Each term in $\mathcal{C}_{\rho}$ scales linearly in $\rho$:
$\frac{\dot{a}^{2}}{\rho}\sim\rho,\quad\rho\dot{f}^{2}\sim\rho,\quad\frac{f^{2}a^{2}}{\rho}\sim\rho,\quad\rho(f^{2}-1)^{2}\sim\rho.$ (81)
Thus $\mathcal{C}_{\rho}\sim\rho$ and $\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha+1}\to 0$ for $\alpha>-1$.
At infinity ($\rho\to\infty$): Both fields approach their vacuum values exponentially,
$a\sim e^{-m_{A}\rho},\quad f-1\sim e^{-m_{H}\rho},$ (82)
where $m_{A}$ and $m_{H}$ are the gauge and Higgs masses. All terms in $\mathcal{C}_{\rho}$ decay exponentially, ensuring convergence for all finite $\alpha$.
Then the integral relations are valid for $\alpha>-1$.
@@ 4.4 Electroweak sphaleron
The bosonic sector of the electroweak Lagrangian in the limit $\theta_{W}\to 0$ is
$\mathcal{L}=-\frac{1}{4}W^{a}_{\mu\nu}W^{a\mu\nu}+|D_{\mu}\Phi|^{2}-\frac{\lambda}{4}(|\Phi|^{2}-v^{2})^{2},$ (83)
where $\Phi$ is the Higgs doublet and $W^{a}_{\mu\nu}$ is the SU(2) field strength.
For the spherically symmetric ansatz with gauge profiles $f_{A}$, $f_{B}$ and Higgs profiles $H$, $K$ MatchevVerner2025, the energy functional becomes
$E_{\rm sph}$ $=\frac{4\pi}{g^{2}}\int_{0}^{\infty}\mathcal{G}_{\rho}\,d\rho,$ (84)
where $\rho$ is scaled by $m_{W}$ and
$\mathcal{G}_{\rho}$ $=(\dot{f}_{A})^{2}+(\dot{f}_{B})^{2}+\frac{(f_{A}^{2}+f_{B}^{2}-1)^{2}}{2\rho^{2}}+2m_{W}^{2}\rho^{2}(\dot{H}^{2}+\dot{K}^{2})$
$\quad+m_{W}^{2}U_{\rm int}+\frac{m_{W}^{2}m_{H}^{2}}{2}\rho^{2}(H^{2}+K^{2}-1)^{2},$ (85)
with the interaction potential $U_{\rm int}=(f_{A}^{2}+f_{B}^{2})(H^{2}+K^{2})+2(f_{A}H+f_{B}K)^{2}$.
The auxiliary quantity is
$\mathcal{C}_{\rho}$ $=(\dot{f}_{A})^{2}+(\dot{f}_{B})^{2}+2m_{W}^{2}\rho^{2}(\dot{H}^{2}+\dot{K}^{2})-\frac{(f_{A}^{2}+f_{B}^{2}-1)^{2}}{2\rho^{2}}-m_{W}^{2}U_{\rm int}$
$\quad-\frac{m_{W}^{2}m_{H}^{2}}{2}\rho^{2}(H^{2}+K^{2}-1)^{2}.$ (86)
Unlike the BPST instanton where conformal invariance rendered the $\alpha=1$ identity trivial, here the Higgs mass $m_{H}$ explicitly breaks scale invariance: the $\rho^{2}$ factors in the Higgs kinetic and vacuum terms introduce a definite length scale.
@@
The sphaleron has four profile functions with specific regularity conditions at the origin and exponential decay at infinity.
At the origin ($\rho\to 0$): The regularity condition $f_{A}^{2}(0)+f_{B}^{2}(0)=1$ (typically $f_{A}(0)=1$, $f_{B}(0)=0$) implies
$f_{A}^{2}+f_{B}^{2}-1\sim\rho^{2},\quad H,K\sim\rho.$ (96)
The gauge kinetic terms $(\dot{f}_{A})^{2}+(\dot{f}_{B})^{2}$ remain $O(1)$ at the origin. The most restrictive integral is $I_{\rm GC}^{\alpha}$:
$\int(f_{A}^{2}+f_{B}^{2}-1)^{2}\rho^{\alpha-3}\,d\rho\sim\int\rho^{4}\cdot\rho^{\alpha-3}\,d\rho=\int\rho^{\alpha+1}\,d\rho,$ (97)
which converges for $\alpha>-2$.
At infinity ($\rho\to\infty$): All fields approach their vacuum values exponentially with characteristic scale $m_{W}^{-1}$, ensuring convergence for all finite $\alpha$.
Then virial relations hold for $\alpha>-2$, with no upper bound.
@@ 5 Skyrmions
The Skyrme model provides an effective field theory of pions whose topological solitons, Skyrmions, model baryons in the large-$N_{c}$ limit of QCD Skyrme1962; AdkinsNappiWitten1983. The Lagrangian is
$\mathcal{L}=\frac{f_{\pi}^{2}}{16}\text{Tr}(\partial_{\mu}U^{\dagger}\partial^{\mu}U)+\frac{1}{32e^{2}}\text{Tr}([U^{\dagger}\partial_{\mu}U,U^{\dagger}\partial_{\nu}U]^{2}),$ (98)
where $U\in\text{SU}(2)$ is the chiral field, $f_{\pi}$ is the pion decay constant, and $e$ is the Skyrme parameter.
The fundamental $B=1$ Skyrmion has a spherically symmetric hedgehog ansatz
$U=\exp(i\bm{\tau}\cdot\hat{\bm{r}}f(\rho))=\cos f+i\bm{\tau}\cdot\hat{\bm{r}}\sin f,$ (99)
where $\bm{\tau}$ are Pauli matrices, $\hat{\bm{r}}=\bm{r}/\rho$, and the profile satisfies $f(0)=\pi$ (baryon number at the core) and $f(\infty)=0$ (vacuum at infinity).
@@
$E=4\pi\int_{0}^{\infty}\mathcal{G}_{\rho}\,d\rho,\qquad\mathcal{G}_{\rho}=\rho^{2}\mathcal{E}(f,\dot{f}),$ (100)
@@ 5.2 Standard Skyrme model
The standard model has $c_{2,0,0}=c_{2,2,0}=1$, giving the reduced functional
$\mathcal{G}_{\rho}=\rho^{2}\dot{f}^{2}+2\sin^{2}f+2\dot{f}^{2}\sin^{2}f+\frac{\sin^{4}f}{\rho^{2}},$ (104)
so that $E=4\pi\int_{0}^{\infty}\mathcal{G}_{\rho}\,d\rho$. The four terms represent: Dirichlet (radial) gradient energy $\rho^{2}\dot{f}^{2}$; sigma-model (angular) gradient energy $2\sin^{2}f$; Skyrme term $2\dot{f}^{2}\sin^{2}f$ mixing radial and angular strains; and centrifugal barrier $\sin^{4}f/\rho^{2}$ that prevents collapse.
The auxiliary quantity is
$\mathcal{C}_{\rho}=\rho^{2}\dot{f}^{2}+2\dot{f}^{2}\sin^{2}f-2\sin^{2}f-\frac{\sin^{4}f}{\rho^{2}}.$ (105)
@@ 5.3 Asymptotic universality
The Skyrmion has power-law decay at infinity (massless pions), yielding a finite $\alpha$ range.
At the origin ($\rho\to 0$): The hedgehog boundary condition $f(0)=\pi$ and regularity require
$f=\pi-\beta\rho+O(\rho^{3}),\quad\dot{f}\sim-\beta,\quad\sin f=\sin(\pi-\beta\rho)\sim\beta\rho.$ (109)
Each term in $\mathcal{C}_{\rho}$ scales as $\rho^{2}$:
$\rho^{2}\dot{f}^{2}\sim\rho^{2},\quad\dot{f}^{2}\sin^{2}f\sim\rho^{2},\quad\sin^{2}f\sim\rho^{2},\quad\frac{\sin^{4}f}{\rho^{2}}\sim\rho^{2}.$ (110)
Thus $\mathcal{C}_{\rho}\sim\rho^{2}$ and $\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha+2}\to 0$ for $\alpha>-2$.
At infinity ($\rho\to\infty$): The massless pion field decays as a power law,
$f\sim\frac{C}{\rho^{2}},\quad\dot{f}\sim-\frac{2C}{\rho^{3}},\quad\sin f\approx f\sim\rho^{-2}.$ (111)
The Dirichlet term dominates: $\rho^{2}\dot{f}^{2}\sim\rho^{-4}$. The angular terms decay faster: $\sin^{2}f\sim\rho^{-4}$, $\dot{f}^{2}\sin^{2}f\sim\rho^{-10}$, $\sin^{4}f/\rho^{2}\sim\rho^{-10}$. Thus $\mathcal{C}_{\rho}\sim\rho^{-4}$ and
$\rho^{\alpha}\mathcal{C}_{\rho}\sim\rho^{\alpha-4}\to 0\quad\text{for }\alpha<4.$ (112)
As a consequence the valid range is $2<\alpha<4$. The finite upper bound reflects the power-law decay of massless pions.
@@ 6 Conclusion
The numerical examples illustrate how the $\alpha$-dependence of errors distinguishes between core and tail inaccuracies. For Nielsen-Olesen vortices, the errors grow at negative $\alpha$: the $\alpha=1$ identity held to $0.0005\%$ while $\alpha=-0.5$ showed a $5.7\%$ discrepancy, exposing core-region errors that the global Derrick test averages away. For the Bounce solutions, the opposite pattern emerged: errors grew monotonically with $\alpha$, from $0.001\%$ at $\alpha=-2$ to $0.03\%$ at $\alpha=2$, indicating that numerical inaccuracies were concentrated in the asymptotic tail where the bounce approaches the false vacuum. The $\alpha$-family thus gives a clear signature: core errors manifest at negative $\alpha$, while tail errors manifest at positive $\alpha$.
@@ Acknowledgements.
The author thanks Manuel Torres Labansat for discussions on vacuum decay and bounce solutions.